# High-Strength Amorphous Silicon Carbide for Nanomechanics

### v1

@@ I Introduction
Advances in nanotechnology have revolutionized a broad spectrum of fields, with the development of tensile-loaded, thin-film mechanical devices playing a pivotal role in state-of-the-art force, acceleration, and displacement sensing [1, 2, 3, 4, 5, 6, 7]. Two approaches are used to boost the sensitivity of nanomechanical resonators under tensile loads. One approach fabricates the resonators using different thin-film materials in pursuit of films with low mechanical loss tangents resulting in higher intrinsic mechanical quality factors. In room temperature environments, high-tensile amorphous silicon nitride (a-SiN) nanomechanical resonators have marked some of the best performing devices in ultra-sensitive mechanical detectors [8, 9, 10, 11, 12, 13]. Despite the fact that crystalline thin film materials (e.g. crystalline silicon (c-Si) [14], crystalline silicon carbide (c-SiC) [15, 16]) and graphene are expected to have higher theoretical limits, their projected performance relies on having perfect crystal structures with no defects (e.g. edge defects). Additionally it is difficult to practically attain crystalline thin-films [15, 14] which can be easily deposited, have good film isotropy [17] and few lattice imperfections [15, 14]. By nanostructuring edges into free-standing crystalline devices, one introduces a form of defect by exposing the edge of the crystal where fracture can initiate under high-tensile loads [18, 19, 20].
@@
The other approach to attain state-of-the-art sensors is by innovating designs of the nanomechanical resonators’ geometries to have high stresses at crucial regions of the resonator when it vibrates. Ultimately the design space of these resonators is constrained by the thin film materials’ tensile fracture limits or ultimate tensile strength (UTS). The UTS of a thin-film is fundamentally lower after being nanostructured into a suspended geometry with edges, since the introduced crystalline defects such as dislocations can facilitate the propagation of cracks [21, 22]. For example, the UTS of a-SiN thin film has been shown to be 6.8 GPa [23]. To date, only crystalline and 2D materials have experimentally demonstrated UTS surpassing 10 GPa after being top-down nanofabricated [24, 25, 26, 19]. Among 2D crystalline materials, graphene harbors one of the highest theoretical UTS [27, 26], but practically reaching the limit is also challenging due to lattice imperfections [28], atomically irregular edges [19], or sparser grain boundaries [29] resulting from nanostructuring processes, which lead to a reduced fracture limit when it is tensile-loaded. In this regard, amorphous thin films with high UTS offer more design freedom for free-standing nanostructures, due to their lack of both crystalline defects and sensitivity to notches [30, 31, 32, 33]. Apart from allowing the enhancement of the Q factor of nanomechanical resonators, higher material UTS can enable the devices to perform better in diverse and harsh environments.
Among the materials processing the highest ultimate tensile strengths [24, 34, 35], silicon carbide (SiC) is attracting rising interest and demand both in industry and in academia in recent years due to its exceptional mechanical, chemical, electrical and optical properties [36, 37, 38, 39, 40, 41, 42]. SiC exists in three forms: single crystalline, poly-crystalline, and amorphous, all of which can be fabricated into thin films for various applications [36]. Among them, amorphous SiC (a-SiC) thin film offers distinct advantages over their crystalline counterparts [43, 44, 45, 46, 47, 48, 19, 25], including lower deposition temperature, compatibility with various substrates, isotropic physical properties, and the ability to deposit on large wafer scales. The widespread applications of a-SiC range from protective coatings to avoid mechanical wear [49] and chemical corrosion [50, 51], low-loss deposited dielectrics [52], device and functional layers for solar cells [53, 54, 55], MEMS sensors [56, 57, 58, 44] and electron transparent windows [43], low-loss optical resonators [59] and integrated photonics [60, 61], bio-molecular [62] and medical [63] applications, as well as gratings with nano-pillars [64, 65], to name a few. The versatility of a-SiC makes it a promising material for the high-yield production of integrated mechanical, electronic and photonic devices on a chip, paving the way for applications in sensing [6, 66], transduction [67], and quantum technology [68, 69].
@@ II Fabrication of amorphous SiC resonators
In pursuit of thin film materials for nanomechanical resonators with low mechanical dissipation, high film quality and high tensile stress are desirable. The Low Pressure Chemical Vapor Deposition (LPCVD) technique is preferred for these requirements, since its low pressure and high temperature deposition environment ensures lower defect density and higher thermal stress. The non-stoichiometric LPCVD a-SiC films used in this paper are deposited with different gas flow ratios (GFR) between SiHCl and 5% CH in H (GFR=2,3,4), various deposition pressures (170 and 600 mTorr), and on both silicon and fused silica substrates (Table 1). This variation of deposition parameters allows us to systematically characterize the mechanical properties of LPCVD a-SiC thin films. All a-SiC thin films were deposited for the same period of time (3 hours 47 minutes) at a temperature of 760C in order to better determine the effect of various deposition environments while maintaining the films in the amorphous form [73]. With the fabrication process demonstrated in the Supporting Information (H), nanomechanical resonators made of a-SiC can be suspended over the substrates with high yield using dry etching processes due to their extremely high chemical selectivity.
@@
Better chemical stability and inertness of the sensing components can significantly improve the sensors’ reliability, particularly for their operation under chemically harsh environments. Meanwhile, thin film materials’ chemical inertness allows them to be deposited on various substrates, patterned, and then suspended as nanomechanical resonators by removing the substrate underneath (i.e. undercutting). A high selectivity between the thin film and the substrate allows for higher yield and accuracy in fabricating suspended nanostructures . Similar to their crystalline counterpart, LPCVD a-SiC thin films have been reported to have very high chemical inertness to various wet etchants [43]. Likewise, we found that a-SiC also has high chemical inertness to the widely used dry etchants, such as SF isotropic plasma etching for silicon substrate, and vapor hydrofluoric acid etching for silicon oxide substrate, as illustrated in Figure 1(a). The excellent chemical stability implies high selectivity between a-SiC and various commonly used substrates during undercutting, as shown in Figure 1(b). Dry etchants are preferred for suspending high-aspect-ratio nanomechanical structures, since they help to avoid stiction during liquid etchant evaporation and thus improve the yield rate of working devices. To take advantage of the superior chemical inertness, we fabricated nanomechanical resonators with continuous films down to 5nm, as shown in SEM pictures in Supporting Information (F).
@@
Moreover, the undercut method based on dry etchants introduces little perturbation to the suspended nano-structure, making it possible to perform tensile testing on chips. In Section IV, we fabricate suspended nano-structures with different maximum tensile stresses to accurately determine the ultimate tensile strength (UTS) of a-SiC films. As a result, we demonstrate that a-SiC films have UTS up to 10-12 GPa, which are the highest among amorphous materials after patterning and are approaching the UTS of strong materials like c-SiC [24] and graphene nano-ribbons [19], both of which are known for their high UTS. The comparison of UTS between LPCVD a-SiC and other materials commonly used for nanomechanics is shown in Figure 1(c).
@@ III Mechanical property characterization with resonance method
In order to design desired nanomechanical resonators with a specific thin film material, it is necessary to accurately characterize the material’s mechanical parameters, such as film stress, Young’s modulus, Poisson ratio and density. Various methods are developed to measure these parameters, including static methods, like nano-indentation [78, 79] and dynamic methods like resonance response [80, 81, 82, 83]. Many studies aiming to design high-performance nanomechanical resonators have relied on mechanical parameter values obtained from literature without considering potential variations of thin film properties due to different deposition environments, such as commonly used materials like a-SiN [10, 12], c-Si [14], and c-SiC [15, 16]. While these adaptations are usually reasonable and align well with experimental results, characterizing the exact parameters of the materials used would be beneficial when exploring the optimal performance of nanomechanical resonators [84, 80]. In this section, we present a simple and universal method to systematically characterize the important mechanical parameters of LPCVD a-SiC thin films.
@@
Recipe |
GFR |
P (mTorr) |
Substrate |
(kg/m) |
UTS (GPa) |
a-SiCR2 |
2 |
600 |
Silicon |
71 |
760 |
2830 |
223 |
0.203 |
5175 |
12.04 |
a-SiCR2FS |
2 |
600 |
Fuse Silica |
71 |
1596 |
2830 |
223 |
0.203 |
4554 |
12.04 |
a-SiC170 |
2 |
170 |
Silicon |
86 |
636 |
2966 |
220 |
0.218 |
4485 |
10.27 |
a-SiCR3 |
3 |
600 |
Silicon |
81 |
960 |
2962 |
210 |
0.199 |
4692 |
11.12 |
a-SiCR4 |
4 |
600 |
Silicon |
137 |
670 |
3087 |
200 |
0.162 |
1035 |
3.5 |
Parameters of a-SiCR2FS are adapted from a-SiCR2 due to its transparent substrate
@@ IV Ultimate Tensile Strength of amorphous SiC
Ultimate tensile strength (UTS), or sometimes coinciding with yield strength for brittle materials such as a-SiC [35], describes the maximum tensile stress a material can endure before breaking while being stretched. Materials with higher UTS not only allows them to operate more reliably as mechanical sensors or coating in harsh environments, but also enlarges the design space for nanomechanical resonators. High UTS has been shown for nanowires fabricated with various materials, whose small cross-section areas minimize the appearance of defects [85, 86], and for nanomechanical membranes without nano-patterning to avoid the presence of rough sidewalls [87]. However, both scenarios above do not allow for further shape modification, reducing interest in their potential for various applications. While the crystalline form of materials usually tend to be mechanically stronger than their amorphous forms due to long-range order, examples such as glassy metal [88] and synthesized AM-III carbon [89] demonstrate extraordinary mechanical properties comparable to their celebrated crystalline counterparts in terms of fracture toughness and yield strength, or hardness and compressive strength, respectively. This correspondence remains between c-SiC and a-SiC. While c-SiC has shown a UTS as high as 12-18 GPa via micro-pillars so far [24], a-SiC nanowires have been measured to have a UTS up to 8.8 GPa via a tensile test with its two ends fixed by silver epoxy [35], which is higher than the ones shown for LPCVD a-SiN (6.8 GPa [23]) and Si (7.6 GPa [70]).
@@
The survival of the suspending hour-glass-shaped device with the tether length below 50 um, corresponds to a UTS above 10 GPa for a-SiCR2. Similarly, we can identify the UTS for all a-SiC thin films used in this study to be higher than 10 GPa, as shown in the histograms of ratios of survival devices in Figure 3(d). The histograms also show that, with relatively higher deposition pressure and lower gas flow ratios, a maximum UTS up to 12 GPa can be achieved with a-SiCR2, which is almost twice that of the UTS shown for nanostructured LPCVD a-SiN films. The measured UTS of a-SiCR4 is below 3.5 GPa, which is not attractive for further characterization. In the future, with a larger number of fabricated devices and a denser range of tether lengths, one can determine the UTS of the LPCVD a-SiC thin films more precisely. In practice, the nanopatterning with electron beam lithography can readily achieve an accuracy of 10 nm, which allows for the method’s accuracy to be as low as 1.2 MPa on a-SiCR2, i.e. an error of less than 0.2% when measuring the UTS. Higher UTS is found for recipes deposited with lower gas flow ratios (a-SiCR2/3/4), which might due to a higher carbon composition in the thin film [73], and C-C chemical bonds are stronger than Si-C and Si-Si bonds [91]. For a-SiC films deposited with different pressure, a-SiCR2 (600 mTorr) is found to have a higher UTS, while a-SiC170 (170 mTorr) exhibits better yield under lower concentrated stresses as shown by the survival rates. According to the relationship between strength and Young’s modulus of SiC shown in [35], UTS (or fracture strength) is 5.3% of , therefore the theoretically predicted UTS for a-SiCR2/a-SiC170/a-SiCR3, are 11.82/11.66/11.13 GPa, respectively, matching well with the experimentally extracted data from the survived devices 12.04/10.27/11.12 GPa shown in Table 1. The small offset for the values of a-SiC170 may be due to its rougher surface as shown in Supporting Information (G).
@@ V Intrinsic Quality Factor and High Q Mechanical Resonators
In this section we characterize the intrinsic quality factor of LPCVD a-SiC, then design and fabricate high-Q nanomechanical resonators with it. High mechanical quality (Q) factor nanomechanical resonators are desirable for various applications, ranging from precise force/acceleration sensing [6, 1], microwave-to-optical conversion [67], to quantum optomechanics [66, 94]. Following the method introduced by LIGO [95], the field of strain engineering is advancing rapidly, boosting the Q factor of nanomechanical resonators by several orders of magnitude. A variety of strategies have been proposed aiming to improve the Q factors of tensile-loaded nanomechanical resonators. These include patterning 2D geometries appropriately [9, 11, 13, 12, 96, 97, 98], modifying mass distribution [99, 2] and mode of interest (e.g., from fundamental to higher order or from flexural to torsional modes [2]), in-situ annealing for surface cleaning [100], as well as cooling down to cryogenic temperatures [101, 14]. The methods mentioned above can benefit from utilizing the LPCVD a-SiC thin film we characterized in this work, due to its high deposition film tensile stress, superior chemical resistivity, and impressive ultimate tensile strength.
@@
The intrinsic quality factors of a-SiC thin films are identified by experimentally measuring the Q factors of phononic crystal (PnC) nanostrings [10], whose many spurious loss mechanisms are eliminated, and dissipation dilution factor is well defined, leading to an expected intrinsic Q factor . For thin nanomechanical resonators, can be assumed to depend linearly on the film thickness, since it is predominantly determined by surface loss rather than bulk loss [92]. We fabricate a series of uniformly corrugated high-aspect-ratio (PnC) nanostrings with a length of 4 mm, varying unit-cell lengths and defect lengths in the middle (Figure 4(a)), leading to PnC nanostrings with unit-cell numbers from 20 to 44. The widths of the wide and narrow parts of the nanostrings are 3 um and 1 um respectively. The vibration amplitude of the nanostrings as a function of frequency is acquired (Figure 4(b)) with a custom balanced homodyne detection interferometer at the vacuum environment of mbar (see Supporting Information (I)). Using the ringdown method, the Q factors of defect modes for each PnC nanostring are measured. For example, the ones of 10 unit-cells PnC nanostrings fabricated wtih a-SiCR2 and a-SiCR2FS are plotted in Figure 4(d). Using FEM simulation, the dilution factor of each PnC nanostring geometry can be numerically calculated. Together with the Q factors of the corresponding nanostring measured experimentally, the intrinsic Q factor of the different a-SiC thin films are determined. For example, the of a-SiCR2 and a-SiCR2FS are shown in Figure 4(c) and (e) respectively. The of the other a-SiC films are shown in Table 1, and the corresponding measurement data can be found in Supporting Information (D). In order to compare the of a-SiC thin films with different thicknesses, we present them with the unit per 100 nm, as of a thin film is shown to be a function of thickness [84]. Deposited with the same recipe, a-SiCR2 (5175/100 nm) and a-SiCR2FS (4554/100 nm) have similar . The similar performance on the transparent substrate allows for integrating high-Q nanomechanical sensors into free-space optical systems in a practical manner. By reasonably assuming the films have similar mechanical properties on different substrates, a-SiCR2FS is measured to have a deposition stress of 1596 MPa, a factor of two higher than a-SiCR2 due to a larger thermal expansion coefficients difference between the a-SiC thin film and fused silica substrate. Worth noting is that the of a-SiCR2 is the highest among all LPCVD a-SiC investigated, indicating that a lower gas flow ratio (GFR=2), i.e., more carbon content [73], and a moderate deposition pressure (600 mTorr) is beneficial to have better film quality.
@@
In addition, the quality factor-frequency product of the optimized LPCVD a-SiC tapered PnC nanostring is , which is significantly higher than the quantum limit . This paves the way towards engineering quantum states in room temperature environments [94, 9]. This high quality factor of the nanomechanical resonator with an effective mass kg corresponds to a force sensitivity of 7.7 aN/Hz at room temperature, which is comparable to a typical atomic force microscope cantilever operating at liquid helium temperature. With the high quality factor shown above, LPCVD a-SiC is shown to be the third material that can reach at room temperature using strain engineering, after conventional a-SiN [10] and strained silicon [14]. Moreover, the superior chemical and mechanical properties of LPCVD a-SiC allow for the fabrication of thinner and stronger resonators, enabling it to be more compatible with the dissipation dilution method. With advantages such as a relatively simple and low-cost fabrication process, compatibility with various substrates, including transparent ones, and its potential to perform better and more stably in harsh environments as high-Q nanomechanical resonators, LPCVD a-SiC is a promising material for fabricating commercial mechanical sensors.

### v2

@@ I Introduction
Advances in nanotechnology have revolutionized various fields, with the development of tensile-loaded, thin-film mechanical devices playing a pivotal role in state-of-the-art force, acceleration, and displacement sensing [1, 2, 3, 4, 5, 6, 7]. Two approaches are used to boost the sensitivity of nanomechanical resonators under tensile loads. One approach fabricates the resonators using different materials in pursuit of films with inherent high-stress and low mechanical loss tangents (i.e high mechanical quality factors). In room temperature environments, high-tensile amorphous silicon nitride (a-Si3N4) nanomechanical resonators have marked some of the best performing devices in ultra-sensitive mechanical detectors [8, 9, 10, 11, 75, 12]. Crystalline thin film materials (e.g. crystalline silicon (c-Si) [13], crystalline silicon carbide (c-SiC) [14, 15]) and graphene are expected to have higher theoretical limits, but their projected performance relies on having perfect crystal structures with no defects (incl. edge defects). Additionally it is difficult to attain crystalline films [14, 13] which can be easily deposited, have good film isotropy [16] and few lattice imperfections [14, 13].
@@
The other approach to boost sensor performance involves innovative resonator designs that concentrate stress in key areas. These designs are constrained by the thin film materials’ tensile fracture limits or ultimate tensile strength (UTS). Nanostructuring reduces the UTS due to introduced crystalline defects [17, 18]. For example, the UTS of a-Si3N4 thin film has been shown to be 6.8 GPa [19]. To date, only crystalline and 2D materials have experimentally demonstrated UTS surpassing 10 GPa after being top-down nanofabricated [20, 21, 22, 23]. Among 2D crystalline materials, graphene harbors one of the highest theoretical UTS [24, 22], but practically reaching the limit is also challenging due to lattice imperfections [25], atomically irregular edges [23], or sparser grain boundaries [26] resulting from nanostructuring processes, which lead to a reduced fracture limit when it is tensile-loaded. In this regard, amorphous thin films with high UTS offer more design freedom for free-standing nanostructures, due to their lack of both crystalline defects and sensitivity to notches [27, 28, 29, 30]. Apart from allowing the enhancement of the Q factor of nanomechanical resonators, higher material UTS can enable the devices to perform better in diverse and harsh vibrational environments.
Amorphous SiC (a-SiC) thin film is gaining traction due to its remarkable mechanical strength and versatile properties [31, 32, 33]. It holds unique advantages over its crystalline counterparts, such as lower deposition temperature and adaptability to various substrates [34, 35], enabling deposition on large wafer scales. This material stands out in applications requiring protective coatings and in the development of MEMS sensors and integrated photonics, due to its resilience to mechanical wear [36] and chemical corrosion [37]. Its potential in high-yield production of diverse devices paves the way for advancements in sensing [6] and quantum technology [38].
@@ II Fabrication of amorphous SiC resonators
In pursuit of thin film materials for nanomechanical resonators with low mechanical dissipation, high film quality and high tensile stress are desirable. The Low Pressure Chemical Vapor Deposition (LPCVD) technique is preferred for these requirements, since its low pressure and high temperature deposition environment ensures lower defect density and higher thermal stress. The non-stoichiometric LPCVD a-SiC films used in this paper are deposited with different gas flow ratios (GFR) between SiH2Cl2 and 5% C2H2 in H2 (GFR=2,3,4), various deposition pressures (170 and 600 mTorr), and on both silicon and fused silica substrates (Table 1). This variation of deposition parameters allows us to systematically characterize the mechanical properties of LPCVD a-SiC thin films. All a-SiC thin films were deposited for the same period of time (3 hours 20 minutes) at a temperature of 760∘C in order to better determine the effect of various deposition environments while maintaining the films in the amorphous form [43]. With the fabrication process demonstrated in the Experimental Section, nanomechanical resonators made of a-SiC can be suspended over the substrates with high yield using dry etching processes due to their extremely high chemical selectivity.
@@
Higher chemical stability and inertness of thin films is particularly useful for suspending nanostructures; usually a crucial and difficult step in fabricating free-standing structures. LPCVD a-SiC films can be deposited on various substrates, patterned, and then suspended as nanomechanical resonators by removing the substrate underneath (i.e. undercutting). A high selectivity between the thin film and the substrate allows for higher yield and accuracy in fabricating suspended nanostructures. Similar to their crystalline counterpart, LPCVD a-SiC thin films have been reported to have very high chemical inertness to various wet etchants [34]. Likewise, we found that a-SiC also has high chemical inertness to the widely used dry etchants, such as SF6 isotropic plasma etching for silicon substrate, and vapor hydrofluoric acid etching for silicon oxide substrate, as illustrated in Figure 1(a). The excellent chemical stability implies high selectivity between a-SiC and various commonly used substrates during undercutting, as shown in Figure 1(b). Dry etchants are preferred for suspending high-aspect-ratio nanomechanical structures, since they help to avoid stiction during liquid etchant evaporation and thus improve the yield rate of working devices. To demonstrate the range of devices possible with higher chemical inertness, we fabricated nanomechanical resonators with continuous films down to 5nm, as shown in SEM pictures in Supporting Information (F). In the Supporting Information (H) we show the measured chemical composition of these films, and demonstrate the chemical integrity of the 5nm SiC after etching.
@@
The undercut method based on dry etchants introduces little perturbation to the suspended nano-structure, making it possible to perform delicate, on-chip tensile testing. The precision of the undercut method is pivotal for the reliability of material strength testing as it mitigates the risks associated with introducing fractures and defects during the loading/gluing of a material into a tensile testing setup, which could compromise the integrity of maximum tensile strength measurements. This chemical inertness of a-SiC serves a functional role in the fabrication process, allowing for a high degree of compatibility with various undercuts and ensuring the integrity of the nanostructures during the suspension process. The inertness contributes to the overall reliability and accuracy of the UTS measurements by inducing minimal forces during both fabrication and testing. In Section IV, we fabricate suspended nano-structures with different maximum tensile stresses to accurately determine the ultimate tensile strength (UTS) of a-SiC films. As a result, we demonstrate that a-SiC films have UTS up to 10-12 GPa, which are the highest among amorphous materials after patterning and are approaching the UTS of strong materials like c-SiC [20] and graphene nano-ribbons [23], both of which are known for their high UTS. The comparison of UTS between LPCVD a-SiC and other materials commonly used for nanomechanics is shown in Figure 1(c).
@@ III Mechanical property characterization with resonance method
In order to design desired nanomechanical resonators with a specific thin film material, it is necessary to accurately characterize the material’s mechanical parameters, such as film stress, Young’s modulus, Poisson ratio and density. Various methods are developed to measure these parameters, including static methods, like nano-indentation [49, 50] and dynamic methods like resonance response [51, 52, 53, 54]. Many studies aiming to design high-performance nanomechanical resonators have relied on mechanical parameter values obtained from literature without considering potential variations of thin film properties due to different deposition environments, such as commonly used materials like a-Si3N4 [10, 75], c-Si [13], and c-SiC [14, 15]. While these adaptations are usually reasonable and align well with experimental results, characterizing the exact parameters of the materials used would be beneficial when exploring the optimal performance of nanomechanical resonators [55, 51]. In this section, we present a simple and universal method to systematically characterize the important mechanical parameters of LPCVD a-SiC thin films.
@@
| Recipe | GFR | P (mTorr) | Substrate | (nm) | (MPa) | (kg/m3) | (GPa) | UTS (GPa) | ||
| a-SiCR2 | 2 | 600 | Silicon | 71 | 760 | 2830 | 223 | 0.203 | 5175 | 12.04 (0.72) |
| a-SiCR2FS | 2 | 600 | Fused Silica | 82 | 1404 | 2555 | 187 | 0.222 | 4416 | - |
| a-SiC170 | 2 | 170 | Silicon | 86 | 636 | 2966 | 220 | 0.218 | 4485 | 10.27 (0.62) |
| a-SiCR3 | 3 | 600 | Silicon | 81 | 960 | 2962 | 210 | 0.199 | 4692 | 11.12 (0.45) |
| a-SiCR4 | 4 | 600 | Silicon | 137 | 670 | 3087 | 200 | 0.162 | 1035 | 3.5 |
@@ IV Ultimate Tensile Strength of amorphous SiC
Ultimate tensile strength (UTS), or sometimes coinciding with yield strength for brittle materials such as a-SiC [48], describes the maximum tensile stress a material can endure before breaking while being stretched. High UTS has been shown for nanowires fabricated with various materials, whose small cross-section areas minimize the appearance of defects [56, 57], and for nanomechanical membranes without nano-patterning to avoid the presence of rough sidewalls [58]. However, both scenarios above do not allow for further shape modification, reducing interest in their potential for various applications. While the crystalline form of materials usually tend to be mechanically stronger than their amorphous forms due to long-range order, examples such as glassy metal [59] and synthesized AM-III carbon [60] demonstrate extraordinary mechanical properties comparable to their celebrated crystalline counterparts in terms of fracture toughness and yield strength, or hardness and compressive strength, respectively. This correspondence remains between c-SiC and a-SiC. While c-SiC has shown a UTS as high as 12-18 GPa via micro-pillars so far [20], a-SiC nanowires have been measured to have a UTS up to 8.8 GPa via a tensile test with its two ends fixed by silver epoxy [48], which is higher than the ones shown for LPCVD a-Si3N4 (6.8 GPa [19]) and Si (7.6 GPa [40]).
@@
The survival of the suspending hour-glass-shaped device with the tether length below 50 um, corresponds to a UTS above 10 GPa for a-SiCR2. Similarly, we can identify the UTS for all a-SiC thin films used in this study to be higher than 10 GPa, as shown in the histograms of ratios of survival devices in Figure 3(d). The histograms also show that, with relatively higher deposition pressure and lower gas flow ratios, a maximum UTS up to 12 GPa can be achieved with a-SiCR2, which is almost twice that of the UTS shown for nanostructured LPCVD a-Si3N4 films. The measured UTS of a-SiCR4 is below 3.5 GPa, which is not attractive for further characterization. In the future, with a larger number of fabricated devices and a denser range of tether lengths, one can determine the UTS of the LPCVD a-SiC thin films more precisely. In practice, the nanopatterning with electron beam lithography can readily achieve an accuracy of 10 nm, which allows for the method’s accuracy to be as low as 1.2 MPa on a-SiCR2, i.e. an error of less than 0.2% when measuring the UTS. Higher UTS is found for recipes deposited with lower gas flow ratios (a-SiCR2/3/4), which might due to a higher carbon composition in the thin film [43], and C-C chemical bonds are stronger than Si-C and Si-Si bonds [62]. For a-SiC films deposited with different pressure, a-SiCR2 (600 mTorr) is found to have a higher UTS, while a-SiC170 (170 mTorr) exhibits better yield under lower concentrated stresses as shown by the survival rates. According to the relationship between strength and Young’s modulus of SiC shown in [48], UTS (or fracture strength) is 5.3% of , therefore the theoretically predicted UTS for a-SiCR2/a-SiC170/a-SiCR3, are 11.82/11.66/11.13 GPa, respectively, matching well with the experimentally extracted data from the survived devices 12.04/10.27/11.12 GPa shown in Table 1. The small offset for the values of a-SiC170 may be due to its rougher surface as shown in Supporting Information (G).
@@ V Intrinsic Quality Factor and High Q Mechanical Resonators
In this section we characterize the intrinsic quality factor of LPCVD a-SiC, then design and fabricate high-Q nanomechanical resonators with it. High mechanical quality (Q) factor nanomechanical resonators are desirable for various applications, ranging from precise force/acceleration sensing [6, 1], microwave-to-optical conversion [65], to quantum optomechanics [66, 67]. Following the method introduced by LIGO [68], the field of strain engineering is advancing rapidly, boosting the Q factor of nanomechanical resonators by several orders of magnitude. A variety of strategies have been proposed aiming to improve the Q factors of tensile-loaded nanomechanical resonators. These include patterning 2D geometries appropriately [9, 11, 12, 75, 69, 70], modifying mass distribution [71, 2] and mode of interest (e.g., from fundamental to higher order or from flexural to torsional modes [2]), in-situ annealing for surface cleaning [72], as well as cooling down to cryogenic temperatures [73, 13]. The methods mentioned above can benefit from utilizing the LPCVD a-SiC thin film we characterized in this work, due to its high deposition film tensile stress, superior chemical resistivity, and impressive ultimate tensile strength.
@@
The intrinsic quality factors of a-SiC thin films are identified by experimentally measuring the Q factors of phononic crystal (PnC) nanostrings [10], whose many spurious loss mechanisms are eliminated, and dissipation dilution factor is well defined, leading to an expected intrinsic Q factor . For thin nanomechanical resonators, can be assumed to depend linearly on the film thickness, since it is predominantly determined by surface loss rather than bulk loss [63]. We fabricate a series of uniformly corrugated high-aspect-ratio (PnC) nanostrings with a length of 4 mm, varying unit-cell lengths and defect lengths in the middle (Figure 4(a)), leading to PnC nanostrings with unit-cell numbers from 20 to 44. The widths of the wide and narrow parts of the nanostrings are 3 um and 1 um respectively. The vibration amplitude of the nanostrings as a function of frequency is acquired (Figure 4(b)) with a custom balanced homodyne detection interferometer at the vacuum environment of mbar (see schematic in Experimental Section). Using the ringdown method, the Q factors of defect modes for each PnC nanostring are measured. For example, the ones of 10 unit-cells PnC nanostrings fabricated wtih a-SiCR2 and a-SiCR2FS are plotted in Figure 4(d). Using FEM simulation, the dilution factor of each PnC nanostring geometry can be numerically calculated. Together with the Q factors of the corresponding nanostring measured experimentally, the intrinsic Q factor of the different a-SiC thin films are determined. For example, the of a-SiCR2 and a-SiCR2FS are shown in Figure 4(c) and (e) respectively. The of the other a-SiC films are shown in Table 1, and the corresponding measurement data can be found in Supporting Information (D). In order to compare the of a-SiC thin films with different thicknesses, we present them with the unit per 100 nm, as of a thin film is shown to be a function of thickness [55]. Deposited with the same recipe, a-SiCR2 (5175/100 nm) and a-SiCR2FS (4554/100 nm) have similar . The similar performance on the transparent substrate allows for integrating high-Q nanomechanical sensors into free-space optical systems in a practical manner. By reasonably assuming the films have similar mechanical properties on different substrates, a-SiCR2FS is measured to have a deposition stress of 1596 MPa, a factor of two higher than a-SiCR2 due to a larger thermal expansion coefficients difference between the a-SiC thin film and fused silica substrate. Worth noting is that the of a-SiCR2 is the highest among all LPCVD a-SiC investigated, indicating that a lower gas flow ratio (GFR=2), i.e., more carbon content [43], and a moderate deposition pressure (600 mTorr) is beneficial to have better film quality.
@@
In addition, the quality factor-frequency product of the optimized LPCVD a-SiC tapered PnC nanostring is , which is significantly higher than the quantum limit . This paves the way towards engineering quantum states in room temperature environments [67, 9]. This high quality factor of the nanomechanical resonator with an effective mass kg corresponds to a force sensitivity of 7.7 aN/Hz1/2 at room temperature, which is comparable to a typical atomic force microscope cantilever operating at liquid helium temperature. With the high quality factor shown above, LPCVD a-SiC is shown to be the third material that can reach at room temperature using strain engineering, after conventional a-Si3N4 [10] and strained silicon [13]. Moreover, the superior chemical and mechanical properties of LPCVD a-SiC allow for the fabrication of thinner and stronger resonators, enabling it to be more compatible with the dissipation dilution method. With advantages such as a relatively simple and low-cost fabrication process, compatibility with various substrates, including transparent ones, its high-Q performance, LPCVD a-SiC is a promising material for fabricating commercial mechanical sensors.